OpenStudy (anonymous):

i need step by step instructions on how to solve the equation 17= -13 - 8x

OpenStudy (anonymous):

add 13 to both sides, then divide both sides by -8

OpenStudy (anonymous):

could you show me

OpenStudy (anonymous):

First we need to isolate -8x, so lets get rid of the thirteen 17 = -13 - 8x lets add 13 to cancel out the -13, but what you do to oneside, you must do to the other: 17 + 13 = -13 - 8x +13 30 = -8x now lets isolate the (x) by dividing both sides by -8 30/-8 = x 15/-4 = x -15/4 = x

OpenStudy (anonymous):

Answer is x =-15/4